# Rise and fall of a multicomponent droplet in a surrounrdfing fluid: simulation study of a bumpy path

### v1

@@ V Conclusion
We have carried out a 2D simulation of the dynamics of a multicomponent droplet released in a water column using the improved “phase_field” model of TrioCFD code. Such a model can solve several equations of Cahn-Hilliard coupled with Navier-Stokes equations to treat coupled mass transfer and hydrodynamic phenomena. To illustrate both the capacity of the model to solve such problem and the interest of fully coupled simulations we have considered the case of a multicomponent droplet rising and sinking in water. The relative importance of advection and diffusion is a key parameter in this phenomenon. Accordingly, we have carried out a study with different mobilities in order to understand how the diffusion layer is affected by the flow. In our case of study where the flow is buoyancy driven, the effects of considering a fully coupled system go far beyond a correction of the parameters of a simplified model. Indeed, the coupling between the flow and the diffusion leads to a significant increase in the magnitude of the upward motion of the droplet before it starts to fall. The analysis of our results allows us to link this phenomenon with a drag effect of the light diffusion layer around the droplet.
@@
Hence in this work we have shown that the use of fully coupled models for the evolution of multiphase flows can be necessary as soon as there is a significant diffusion from one phase to the other. This is typically the case in many multicomponent systems. This work can find applications in many other fields of application ranging from the study of phase inversion of the corium pool and CO2 sequestration in deep aquifers to microfluidic devices.
@@ Acknowledgements.
AIP Publishing believes that all datasets underlying the conclusions of the paper should be available to readers. Authors are encouraged to deposit their datasets in publicly available repositories or present them in the main manuscript. All research articles must include a data availability statement stating where the data can be found. In this section, authors should add the respective statement from the chart below based on the availability of data in their paper.

### v2

@@ VII Conclusion
The numerical simulation results show that the relative importance of advection and diffusion is a key parameter in the situation considered here. Accordingly, we have carried out a study with different mobilities in order to understand how the diffusion layer is affected by the flow. In our case of study where the flow is buoyancy driven, the effects of considering a fully coupled system go far beyond a correction of the parameters of a simplified model. Indeed, the coupling between the flow and the diffusion leads to a significant increase in the magnitude of the upward motion of the droplet before it starts to fall. The analysis of our results allows us to link this phenomenon with a drag effect of the light diffusion layer around the droplet.
@@
Hence in this work we have shown that the use of fully coupled models for the evolution of multiphase flows can be necessary as soon as there is a significant diffusion from one phase to the other. This is typically the case in many multicomponent systems. This work can find applications in many other fields of application ranging from the study of phase inversion of the corium pool and CO2 sequestration in deep aquifers to microfluidic devices and many other Lohse and Zhang (2020). In addition, the model studied here can be used in a 3D code without any change.
@@ Acknowledgements.
Data will be made available on https://entrepot.recherche.data.gouv.fr/dataverse/cnrs